A uniform solid ball has a few turns of light string wound around it. If the end of the string


is held steady and the ball is allowed to fall under gravity,what is the acceleration of the


center of the ball? (Assume the string remains vertical.)

1
Expert's answer
2021-12-13T10:34:25-0500

"a=g\arcsin{\sin\frac 23}=6.5~\frac m{s^2}."